The patient is a 7 year old child who weighs 52 lbs. Ordered: Rantidine 30 mg...

  1. The patient is a 7 year old child who weighs 52 lbs.

Ordered: Rantidine 30 mg IV q8h.

On hand Rantidine 50 mg/ 2 mL

Pediatric Use:

While limited data exists on the administration of IV rantidine to children, the recommended dose in pediatric patients is for a total daily dose of 2-4 mg/kg, to be divided and administered every 6 – 8 hours. Up to a maximum of 50 mg given every 6 – 8 hours. This recommendation is derived from adult clinical studies and pharmacokinetic data in pediatric patients.

Package insert states:

  1. Is the dosage ordered safe? Yes______ NO______ Show calculations that proves your choice.
  2. How will the medication be given?  
  3. How often will the medication be given?
  4. How much in mL of the medication to be given at prescribed time
  5. What do you do if the dose is not a safe dose?
Answer #1

ANSWER.

A) Calculation of safe dose.

         Ordered dose = 30 mg q8h.

       Recommended dose = 2mg-4 mg/kg.

   Weight of the patient = 52 lbs.

    Convert lbs to kg for calculation

                  1 lbs = 0.45 kg.

So

       Weight of the patient = 52 * 0.45

                                         = 23.6 kg.

   Dose required for this patient = 23.6 * 2 - 23.6 * 4

                                                   = 47.2 mg -94.4 mg.

         Ordered dose = 30 mg q8h (three times a day)

                                = 30 * 3

                                = 90 mg / day.

* Ordered dose (30 mg q8h) is safe for this patient.

B) This medication is given through intravenous route.

C) This medication is given three times a day (q8h) as per order.

D) Calculation of amount(ml) of drug required.

            Formula = Desired dose /Dose in hand * Quantity.

       Here

          Desired dose(D) = 30 mg.

          Dose in hand(H) = 50 mg.

               Quantity = 2ml.

As per above formula

     Amount(ml) of drug to be given = D/H * Q

                                                        = 30/50 * 2

                                                        = 0.6 * 2

                                                        = 1.2 ml.

D) If the dose is not a safe dose withhold the dose and inform the physician to

     change the dosage.